the demand function Q(P) and cost functions C(Q) of a company's are give by the equations:

Q=12000-60P

C(Q0=10000-4Q

where P and Q are the price and quantity, respectively.

What is the company's profit function?


1
Expert's answer
2022-04-18T01:39:09-0400

The profit function is Pr(P) = Q*P - C(P) = Q*P - (10000 - 4Q) = Q(P + 4) - 10000 =

(12000 - 60P)(P + 4) - 10000 = 38000 + 1760P -60P2